# A Survey Analyzing Generalization in Deep Reinforcement Learning

### 04 Jan 2024 v1

@@ 3 How to Achieve Generalization?
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,P,r,\rho_{0},\gamma)$, a policy transforming generalization method $\mathcal{G}_{\pi}$ is given by a sequence of functions $F_{t}:(S\times A\times S\times\mathbb{R})^{t}\to S\times A$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the policy by which $\mathcal{A}$ chooses the next state and action to be $(\tilde{s_{t}},\tilde{a_{t}})=F_{t-1}(H_{t-1})$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the perturbed policy.
All the definitions so far categorize methods to modify training algorithms in order to achieve generalization. However, many such methods for modifying training algorithms have a corresponding method which can be used to test the generalization capabilities of a trained policy. Our final definition captures this correspondence.
@@ 4 Roots of Overestimation in Deep Reinforcement Learning
Many reinforcement learning algorithms compute estimates for the state-action values in an MDP. Because these estimates are usually based on a stochastic interaction with the MDP, computing accurate estimates that correctly generalize to further interactions is one of the most fundamental tasks in reinforcement learning. A major challenge in this area has been the tendency of many classes of reinforcement learning algorithms to consistently overestimate state-action values. Initially the overestimation bias for $Q$-learning is discussed and theoretically justified by Thrun & Schwartz (1993) as a biproduct of using function approximators for state-action value estimates. Following this initial discussion it has been shown that several parts of the deep reinforcement learning process can cause overestimation bias. Learning overestimated state-action values can be caused by statistical bias of utilizing a single max operator (van Hasselt, 2010), coupling between value function and the optimal policy (Raileanu & Fergus, 2021; Cobbe et al., 2021), or caused by the accumulated function approximation error (Boyan & Moore, 1994).
Several methods have been proposed to target overestimation bias for value iteration algorithms. In particular, to solve this overestimation bias introduced by the max operator (van Hasselt, 2010) proposed to utilize a double estimator for the state-action value estimates. Later, the authors also created a version of this algorithm that can solve high dimensional state space problems (Hasselt et al., 2016). Some of the work on this line of research targeting overestimation bias for value iteration algorithms is based on simply averaging the state-action values with previously learned state-action value estimates during training time (Anschel et al., 2017).
While overestimation bias was demonstrated to be a problem and discussed over a long period of time (Thrun & Schwartz, 1993; van Hasselt, 2010), recent studies also further demonstrated that actor critic algorithms also suffer from this issue (Fujimoto et al., 2018).
@@
Citation Proposed Method Environment Reinforcement Learning Algorithm
Mnih et al. (2015) $\epsilon$-greedy ALE DQN
Bellemare et al. (2016) Count-based ALE A3C and DQN
Osband et al. (2016b) RLSVI Tetris Tabular $Q$-learning
Osband et al. (2016a) Bootstrapped DQN ALE DQN
Houthooft et al. (2017) VIME DeepMind Control Suite TRPO
Fortunato et al. (2018) NoisyNet ALE A3C and DQN
Lee et al. (2021) SUNRISE DCS [^1] DeepMind Control Suiteand Atari Soft Actor-Critic and Rainbow DQN
5 The Role of Exploration in Overfitting
The fundamental trade-off of exploration vs exploitation is the dilemma that the agent can try to take actions to move towards more unexplored states by sacrificing the current immediate rewards. While there is a significant body of studies on provably efficient exploration strategies the results from these studies do not necessarily directly transfer to the high dimensional state or action MDPs. The most prominent indication of this is that, even though it is possible to use deep neural networks as function approximators for large state spaces, the agent will simply not be able to explore the full state space. The fact that the agent is able to only explore a portion of the state space simply creates a bias in the learnt value function (Baird, 1995).
@@ 5 The Role of Exploration in Overfitting
In this section, we will go through several exploration strategies in deep reinforcement learning and how they affect policy overfitting. A quite simple version of this is based on adding noise in action selection during training e.g. $\epsilon$-greedy exploration. Note that this is an example of a policy transforming generalization method $\mathcal{G}_{\pi}$ in Definition 3.5 in Section 3. While $\epsilon$-greedy exploration is widely used in deep reinforcement learning (Wang et al., 2016; Hamrick et al., 2020; Kapturowski et al., 2023), it has also been proven that to explore the state space these algorithms may take exponentially long (Kakade, 2003). Several others focused on randomizing different components of the reinforcement learning training algorithms. In particular, (Osband et al., 2016b) proposes the randomized least squared value iteration algorithm to explore more efficiently in order to increase generalization in reinforcement learning for linearly parametrized value functions. This is achieved by simply adding Gaussian noise as a function of state visitation frequencies to the training dataset. Later, the authors also propose the bootstrapped DQN algorithm (i.e. adding temporally correlated noise) to increase generalization with non-linear function approximation (Osband et al., 2016a).
@@
Houthooft et al. (2017) proposed an exploration technique centered around maximizing the information gain on the agent’s belief of the environment dynamics. In practice, the authors use Bayesian neural networks for effectively exploring high dimensional action space MDPs. Following this line of work on increasing efficiency during exploration Fortunato et al. (2018) proposes to add parametric noise to the deep reinforcement learning policy weights in high dimensional state MDPs. While several methods focused on ensemble state-action value function learning (Osband et al., 2016a), Lee et al. (2021) proposed reweighting target Q-values from an ensemble of policies (i.e. weighted Bellman backups) combined with highest upper-confidence bound action selection. Another line of research in exploration strategies focused on count-based methods that use the direct count of state visitations. In this line of work, Bellemare et al. (2016) tried to lay out the relationship between count based methods and intrinsic motivation, and used count-based methods for high dimensional state MDPs (i.e. Arcade Learning Environment). Yet it is worthwhile to note that most of the current deep reinforcement learning algorithms use very simple exploration techniques such as $\epsilon$-greedy which is based on taking the action maximizing the state-action value function with probability $1-\epsilon$ and taking a random action with probability $\epsilon$ (Mnih et al., 2015; Hasselt et al., 2016; Wang et al., 2016; Hamrick et al., 2020; Kapturowski et al., 2023).
@@ 6 Regularization
In this section we will focus on different regularization techniques employed to increase generalization in deep reinforcement learning policies. We will go through these works by categorizing each of them under data augmentation, adversarial training, and direct function regularization. Under each category we will connect these different line of approaches to increase generalization in deep reinforcement learning to the settings we defined in Section 3.
@@
Citation Proposed Method Environment Reinforcement Learning Algorithm
Yarats et al. (2021) DrQ DeepMind Control Suite, ALE DQN
Laskin et al. (2020b) CuRL DeepMind Control Suite, ALE Soft Actor Critic and DQN
Laskin et al. (2020a) RAD DeepMind Control Suite, ProcGen Soft Actor Critic and PPO
Wang et al. (2020) Mixreg ProcGen DQN and PPO
@@ 6.1 Data Augmentation
Several studies focus on diversifying the observations of the deep reinforcement learning policy to increase generalization capabilities. A line of research in this regard focused on simply employing versions of data augmentation techniques (Laskin et al., 2020a; b; Yarats et al., 2021) for high dimensional state representation environments. In particular, these studies involve simple techniques such as cropping, rotating or shifting the state observations during training time. While this line of work got considerable attention, a quite recent study (Agarwal et al., 2021b) demonstrated that when the number of random seeds is increased to one hundred the relative performance achieved and reported in the original papers of (Laskin et al., 2020b; Yarats et al., 2021) on data augmentation training in deep reinforcement learning decreases to a level that might be significant to mention.
@@
While some of the work on this line of research simply focuses on using a set of data augmentation methods (Laskin et al., 2020a; b; Yarats et al., 2021), other work focuses on proposing new environments to train in (Cobbe et al., 2020). The studies on designing new environments to train deep reinforcement learning policies basically aim to provide high variation in the observed environment such as changing background colors and changing object shapes in ways that are meaningful in the game, in order to increase test time generalization. In the line of robustness and test time performance, a more recent work that is also mentioned in Section 6.3 demonstrated that imperceptible data augmentations can cause significant damage on the policy performance and certified robust deep reinforcement learning policies are more vulnerable to these imperceptible augmentations (Korkmaz, 2023).
@@
Within this category some work focuses on producing more observations by simply blending in (e.g. creating a mixture state from multiple different observations) several observations to increase generalization (Wang et al., 2020). While most of the studies trying to increase generalization by data augmentation techniques are primarily conducted in the DeepMind Control Suite or the Arcade Learning Environment (ALE) (Bellemare et al., 2013), some small fraction of these studies (Wang et al., 2020) are conducted in relatively recently designed training environments like ProcGen (Cobbe et al., 2020). Cobbe et al. (2019) focuses on decoupling the training and testing set for reinforcement learning via simply proposing a new game environment CoinRun.
@@
Table 3: Environment and algorithm details for different direct function regularization strategies for trying to overcome overfitting problems in reinforcement learning. Note that most of the methods based on direct function regularization are a form of policy perturbation method $\mathcal{G}_{\pi}$ to overcome overfitting as described in Section 3.
Citation Proposed Method Environment Reinforcement Learning Algorithm
Igl et al. (2019) SNI and IBAC GridWorld and CoinRun Proximal Policy Optimization
Vieillard et al. (2020b) Munchausen RL Atari DQN and IQN
Lee et al. (2020) Network Randomization 2D CoinRun and 3D DeepMind Lab Proximal Policy Optimization
@@
Amit et al. (2020) Discount Regularization GridWorld and Mujoco [^2] Low dimensional setting of Mujoco is used for this study. Twin Delayed DDPG (TD3)
@@
Agarwal et al. (2021a) PSM DDMC and Rectangle Game [^3] Rectangle game is a simple video game with only two actions, ”Right” and ”Jump”. The game has black background and two rectangles where the goal of the game is to avoid white obstacles and reach to the right side of the screen. Agarwal et al. (2021a) is the only paper we encountered experimenting with this particular game. DrQ
Liu et al. (2021) BN and dropout and $L_{2}/L_{1}$ Mujoco PPO, TRPO, SAC, A2C
@@ 6.2 Direct Function Regularization
While some of the work we have discussed so far focuses on regularizing the data (i.e. state observations) as in Section 6.1, some focuses on directly regularizing the function learned with the intention of simulating techniques from deep neural network regularization like batch normalization and dropout (Igl et al., 2019). While some studies have attempted to simulate these known techniques in reinforcement learning, some focus on directly applying them to overcome overfitting. In this line of research, Liu et al. (2021) proposes to use known techniques from deep neural network regularization to apply in continous control deep reinforcement learning training. In particular, these techniques are batch normalization (BN) (Ioffe & Szegedy, 2015), weight clipping, dropout, entropy and $L_{2}/L_{1}$ weight regularization.
@@
Lee et al. (2020) proposes to utilize a random network to randomize the input observations to increase generalization skills of deep reinforcement learning policies, and tests the proposal in the 2D CoinRun game proposed by Cobbe et al. (2019) and 3D DeepMind Lab. In particular, the authors essentially introduce a random convolutional layer to perturb the state observations. Hence, this study is also a clear example of a state transformation generalization method $\mathcal{G}_{S}$ described in Definition 3.3. While this is another example of random state perturbation methods we will further explain in Section 6.3 the worst-case perturbation methods to target generalization in reinforcement learning policies.
@@
Some work employs contrastive representation learning to learn deep reinforcement learning policies from state observations that are close to each other (Agarwal et al., 2021a). The authors of this study leverage the temporal aspect of reinforcement learning and propose a policy similarity metric. The main goal of the paper is to lay out the sequential structure and utilize representation learning to learn generalizable abstractions from state representations. One drawback of this study is that most of the experimental study is conducted in a non-baseline environment (Rectangle game). Even though the authors show surprising results for this particular game, it is not directly indicated that the proposed method would work for high dimensional state representation MDPs such as the Arcade Learning Environment. Malik et al. (2021) studies query complexity of reinforcement learning policies that can generalize to multiple environments. The authors of this study focus on an example of the transition probability transformation setting $\mathcal{G}_{\mathcal{P}}$ in Definition 3.4, and the reward function transformation setting $\mathcal{G}_{R}$ in Definition 3.2.
@@
Another line of study in direct function generalization investigates the relationship between reduced discount factor and adding an $\ell_{2}$-regularization term to the loss function (i.e. weight decay) (Amit et al., 2020). The authors in this work demonstrate the explicit connection between reducing the discount factor and adding an $\ell_{2}$-regularizer to the value function for temporal difference learning. In particular, this study demonstrates that adding an $\ell_{2}$-regularization term to the loss function is equal to training with a lower discount term, which the authors refer to as discount regularization. The results of this study however are based on experiments from tabular reinforcement learning, and the low dimensional setting of the Mujoco environment.
@@
On the reward transformation for generalization setting $\mathcal{G}_{R}$ defined in Definition 3.2, Vieillard et al. (2020b) adds the scaled log policy to the current rewards. To overcome overfitting some work tries to learn explicit or implicit similarity between the states to obtain a reasonable policy (Lan et al., 2021). In particular, the authors in this work try to unify the state space representations by providing a taxonomy of metrics in reinforcement learning. Several studies proposed different ways to include Kullback-Leibler (KL) divergence between the current policy and the pre-updated policy to add as a regularization term in the reinforcement learning objective (Schulman et al., 2015). Recently, some studies argued that utilizing Kullback-Leibler regularization implicitly averages the state-action value estimates (Vieillard et al., 2020a).

### 30 Oct 2024 v2

@@ 3.7 Generalization Through Policy
Let $\mathcal{A}$ be a training algorithm that takes as input an MDP and outputs a policy. Given an MDP $\mathcal{M}=(S,A,\mathcal{P},r,\rho_{0},\gamma)$, a policy transforming generalization method $\mathcal{G}_{\pi}$ is given by a sequence of functions $F_{t}:(\Pi\times S\times A\times S\times\mathbb{R})^{t}\times S\times\Delta(A)\to\Delta(A)$. The method attempts to achieve generalization by running $\mathcal{A}$ on MDP $\mathcal{M}$, but modifying the current policy by which $\mathcal{A}$ chooses the action at time $t$ to be $\hat{\pi_{t}}(s_{t},\cdot)=F_{t-1}(H_{t-1},s_{t},\pi_{t}(s_{t},\cdot))$, where $H_{t-1}$ is the history of algorithm $\mathcal{A}$ when running with the transformed policy.
3.8 Assessing Generalization
All the definitions so far categorize methods to modify either training algorithms and/or the MDP, i.e. learning environment, training data, in order to achieve generalization. However, many such methods for modifying training algorithms have a corresponding method which can be used to assess the generalization capabilities of a trained policy. Our final definition captures this correspondence.
@@ 4 Roots of Overestimation in Deep Reinforcement Learning
Many reinforcement learning algorithms compute estimates for the state-action values in an MDP. Because these estimates are usually based on a stochastic interaction with the MDP, computing accurate estimates that correctly generalize to further interactions is one of the most fundamental tasks in reinforcement learning. A major challenge in this area has been the tendency of many classes of reinforcement learning algorithms to consistently overestimate state-action values. Initially the overestimation bias for $Q$-learning is discussed and theoretically justified by Thrun and Schwartz, (1993) as a biproduct of using function approximators for state-action value estimates. In particular, Thrun and Schwartz, (1993) proves that if the reinforcement learning policy overestimates the state-action values by $\gamma c$ during learning then the Q-learning algorithm will fail to learn optimal policy if $\gamma>\dfrac{1}{1+c}$.
Following this initial discussion it has been shown that several parts of the deep reinforcement learning process can cause overestimation bias. Learning overestimated state-action values can be caused by statistical bias of utilizing a single max operator (van Hasselt,, 2010), coupling between value function and the optimal policy (Raileanu and Fergus,, 2021; Cobbe et al.,, 2021), or caused by the accumulated function approximation error (Boyan and Moore,, 1994).
Several methods have been proposed to target overestimation bias for value iteration algorithms. In particular, van Hasselt, (2010) demonstrated that the expectation of a maximum of a random variable is not equal to maximum of the expectation of a random variable.
$\mathbb{E}[\max_{i}[X_{i}]]\neq\max_{i}[\mathbb{E}[X_{i}]]\>\>\textrm{where}\>\>X=\{X_{1},X_{2},\dots,X_{N}\}$
This clear distinction shows that simple Q-learning is a biased estimator, and to solve this overestimation bias introduced by the max operator van Hasselt, (2010) proposed to utilize a double estimator for the state-action value estimates. In particular, the double estimator for double Q-learning works as follows
$Q^{\textrm{I}}(s,a)\leftarrow Q^{\textrm{I}}(s,a)+\alpha(s,a)(r(s,a)+\gamma Q^{\textrm{II}}(s^{\prime},\max_{a}Q^{\textrm{I}}(s^{\prime},a))-Q^{\textrm{I}}(s^{\prime},a))$
and
$Q^{\textrm{II}}(s,a)\leftarrow Q^{\textrm{II}}(s,a)+\alpha(s,a)(r(s,a)+\gamma Q^{\textrm{I}}(s^{\prime},\max_{a}Q^{\textrm{II}}(s^{\prime},a))-Q^{\textrm{II}}(s^{\prime},a)).$
Later, the authors also created a version of this algorithm that can solve high dimensional state space problems (Hasselt et al.,, 2016). Some of the work on this line of research targeting overestimation bias for value iteration algorithms is based on simply averaging the state-action values with previously learned state-action value estimates during training time (Anschel et al.,, 2017). While overestimation bias was demonstrated to be a problem and discussed over a long period of time (Thrun and Schwartz,, 1993; van Hasselt,, 2010), recent studies also further demonstrated that actor critic algorithms also suffer from this issue (Fujimoto et al.,, 2018).
5 The Role of Exploration in Overfitting
The fundamental trade-off of exploration vs exploitation is the dilemma that the agent can try to take actions to move towards more unexplored states by sacrificing the current immediate rewards. While there is a significant body of studies on provably efficient exploration strategies the results from these studies do not necessarily directly transfer to the high dimensional state or action MDPs. The most prominent indication of this is that, even though it is possible to use deep neural networks as function approximators for large state spaces, the agent will simply not be able to explore the full state space. The fact that the agent is able to only explore a portion of the state space simply creates a bias in the learnt value function (III,, 1995).
@@ 5 The Role of Exploration in Overfitting
Citation Method Learning Environment Algorithm
Mnih et al., (2015) $\epsilon$-greedy Arcade Learning Environment DQN
Bellemare et al., (2016) Count-based Arcade Learning Environment A3C and DQN
Osband et al., 2016b RLSVI Tetris Tabular $Q$
Osband et al., 2016a Bootstrapped DQN Arcade Learning Environment DQN
Houthooft et al., (2016) VIME DeepMind Control Suite TRPO
Fortunato et al., (2018) NoisyNet Arcade Learning Environment A3C and DQN
Lee et al., (2021) SUNRISE DCS [^1] DeepMind Control Suite& ALE SAC & RDQN
Mahankali et al., (2024) Random Latent Arcade Learning Environment PPO
@@
In this section, we will go through several exploration strategies in deep reinforcement learning and how they affect policy overfitting. A quite simple version of this is based on adding noise in action selection during training e.g. $\epsilon$-greedy exploration. Note that this is an example of a policy transforming generalization method $\mathcal{G}_{\pi}$ in Definition 3.7 in Section 3. While $\epsilon$-greedy exploration is widely used in deep reinforcement learning (Wang et al.,, 2016; Hamrick et al.,, 2020; Kapturowski et al.,, 2023), it has also been proven that to explore the state space these algorithms may take exponentially long (Kakade,, 2003). Several others focused on randomizing different components of the reinforcement learning training algorithms. In particular, Osband et al., 2016b proposes the randomized least squared value iteration algorithm to explore more efficiently in order to increase generalization in reinforcement learning for linearly parametrized value functions. This is achieved by simply adding Gaussian noise as a function of state visitation frequencies to the training dataset. Later, the authors also propose the bootstrapped DQN algorithm (i.e. adding temporally correlated noise) to increase generalization with non-linear function approximation Osband et al., 2016a . Recently, Mahankali et al., (2024) proposed to randomize the reward function to enhance exploration in high dimensional observation MDPs where policy gradient algorithms are used to explore. This study is also a clear example of the generalization through rewards as has been explained in Definition 3.4 in Section 3.
@@
Houthooft et al., (2016) proposed an exploration technique centered around maximizing the information gain on the agent’s belief of the environment dynamics. In practice, the authors use Bayesian neural networks for effectively exploring high dimensional action space MDPs. Following this line of work on increasing efficiency during exploration Fortunato et al., (2018) proposes to add parametric noise to the deep reinforcement learning policy weights in high dimensional state MDPs. While several methods focused on ensemble state-action value function learning (Osband et al., 2016a, ), Lee et al., (2021) proposed reweighting target Q-values from an ensemble of policies (i.e. weighted Bellman backups) combined with highest upper-confidence bound action selection. Another line of research in exploration strategies focused on count-based methods that use the direct count of state visitations. In this line of work, Bellemare et al., (2016) tried to lay out the relationship between count based methods and intrinsic motivation, and used count-based methods for high dimensional state MDPs (i.e. Arcade Learning Environment). Yet it is worthwhile to note that most of the current deep reinforcement learning algorithms use very simple exploration techniques such as $\epsilon$-greedy which is based on taking the action maximizing the state-action value function with probability $1-\epsilon$ and taking a random action with probability $\epsilon$ (Mnih et al.,, 2015; Hasselt et al.,, 2016; Wang et al.,, 2016; Hamrick et al.,, 2020; Kapturowski et al.,, 2023).
@@ 6 Regularization
In this section we will focus on different regularization techniques employed to increase generalization in deep reinforcement learning policies. We will go through these works by categorizing each of them under data augmentation, adversarial training, and direct function regularization. Under each category we will connect these different lines of approach to increase generalization in deep reinforcement learning to the settings we defined in Section 3.
@@
Citation Method Environment Algorithm
Yarats et al., (2021) DrQ DCS, Arcade Learning Environment DQN
Laskin et al., 2020b CuRL DCS, Arcade Learning Environment SAC and DQN
Laskin et al., 2020a RAD DeepMind Control Suite, ProcGen SAC and PPO
Korkmaz, (2023) Semantic Changes Arcade Learning Environment DDQN & A3C
Wang et al., (2020) Mixreg ProcGen DQN and PPO
@@ 6.1 Data Augmentation
Several studies focus on diversifying the observations of the deep reinforcement learning policy to increase generalization capabilities. A line of research in this regard focused on simply employing versions of data augmentation techniques (Laskin et al., 2020a, ; Laskin et al., 2020b, ; Yarats et al.,, 2021) for high dimensional state representation environments. In particular, these studies involve simple techniques such as cropping, rotating or shifting the state observations during training time. While this line of work got considerable attention, a quite recent study Agarwal et al., 2021b demonstrated that when the number of random seeds is increased to one hundred the relative performance achieved and reported in the original papers of (Laskin et al., 2020b, ; Yarats et al.,, 2021) on data augmentation training in deep reinforcement learning decreases to a level that might be significant to mention.
@@
While some of the work on this line of research simply focuses on using a set of data augmentation methods (Laskin et al., 2020a, ; Laskin et al., 2020b, ; Yarats et al.,, 2021), other work focuses on proposing new environments to train in (Cobbe et al.,, 2020). The studies on designing new environments to train deep reinforcement learning policies basically aim to provide high variation in the observed environment such as changing background colors and changing object shapes in ways that are meaningful in the game, in order to increase test time generalization. In the line of robustness and test time performance, a more recent work that is also mentioned in Section 6.3 demonstrated that imperceptible semantically meaningful data augmentations can cause significant damage on the policy performance and certified robust deep reinforcement learning policies are more vulnerable to these imperceptible augmentations (Korkmaz, 2021a, ; Korkmaz,, 2023).
@@
Within this category some work focuses on producing more observations by simply blending in (e.g. creating a mixture state from multiple different observations) several observations to increase generalization (Wang et al.,, 2020). While most of the studies trying to increase generalization by data augmentation techniques are primarily conducted in the DeepMind Control Suite or the Arcade Learning Environment (ALE) (Bellemare et al.,, 2013), some small fraction of these studies (Wang et al.,, 2020) are conducted in relatively recently designed training environments like ProcGen (Cobbe et al.,, 2020). In the line of research proposing learning environments Dennis et al., (2020) proposed unsupervised environment design by changing the environment parameters to asses generalization for maze structured environments by minimax training where the ”adversary” creating an environment for the policy to solve a task with goal and obstacles as an underspecified parameter. Cobbe et al., (2019) focuses on decoupling the training and testing set for reinforcement learning via simply proposing a new game environment CoinRun.
@@
Table 3: Environment and algorithm details for different direct function regularization strategies for trying to overcome overfitting problems in reinforcement learning. Note that most of the methods based on direct function regularization are a form of algorithmic generalization $\mathcal{G}_{\mathbb{A}}$ to overcome overfitting as described in Section 3.
Citation Proposed Method Learning Environment
Igl et al., (2019) SNI and IBAC GridWorld and CoinRun
Vieillard et al., 2020b Munchausen RL Arcade Learning Environment
Lee et al., (2020) Network Randomization 2D CoinRun and 3D DeepMind Lab
@@
Amit et al., (2020) Discount Regularization GridWorld and MuJoCo [^2] Low dimensional setting of MuJoCo is used for this study (Todorov et al.,, 2012).
@@
Agarwal et al., 2021a PSM DDMC and Rectangle Game [^3] Rectangle game is a simple video game with only two actions, ”Right” and ”Jump”. The game has black background and two rectangles where the goal of the game is to avoid white obstacles and reach to the right side of the screen. Agarwal et al., 2021a is the only paper we encountered experimenting with this particular game.
Liu et al., (2021) BN and dropout and $L_{2}/L_{1}$ MuJoCo
@@ 6.2 Direct Function Regularization
While some of the work we have discussed so far focuses on regularizing the data (i.e. state observations) as in Section 6.1, some focuses on directly regularizing the function learned with the intention of simulating techniques from deep neural network regularization like batch normalization and dropout (Igl et al.,, 2019). While some studies have attempted to simulate these known techniques in reinforcement learning, some focus on directly applying them to overcome overfitting. In this line of research, Liu et al., (2021) proposes to use known techniques from deep neural network regularization to apply in continuous control deep reinforcement learning training. In particular, these techniques are batch normalization (BN) (Ioffe and Szegedy,, 2015), weight clipping, dropout, entropy and $L_{2}/L_{1}$ weight regularization. All these methods fall under the algorithmic generalization category $\mathcal{G}_{\mathbb{A}}$ as described in Section3.
@@
Lee et al., (2020) proposes to utilize a random network to essentially achieve a version of randomization in the input observations to increase generalization skills of deep reinforcement learning policies, and tests the proposal in the 2D CoinRun game proposed by Cobbe et al., (2019) and 3D DeepMind Lab (deepmindlab). In particular, the authors essentially introduce a random convolutional layer to achieve this objective. This study is an example of an algorithmic generalization method $\mathcal{G}_{\mathbb{A}}$ described in Definition 3.3 when the single layer random network is not placed at the first layer of the deep neural network. However, when this single layer random network is placed at the first layer of the neural network, this method is essentially just introducing some noise to the state observations of the policy, thus this is an example of state transforming generalization. When this single random layer is placed other than first, the method is no longer a state transforming generalization method because the states are not modified before they have been observed by the algorithm, but rather implicitly changed due to a random convolutional layer added in the architecture. We will further provide clear instances of the state transformation generalization also in Section 6.3 when the worst-case perturbation methods to target generalization in reinforcement learning policies are explained.
Table 4: Algorithm details for different direct function regularization strategies for trying to overcome overfitting problems in reinforcement learning. Note that most of the methods based on direct function regularization are a form of algorithmic generalization $\mathcal{G}_{\mathbb{A}}$ to overcome overfitting as described in Section 3.
Citation Proposed Method Reinforcement Learning Algorithm
Igl et al., (2019) SNI and IBAC Proximal Policy Optimization (PPO)
Vieillard et al., 2020b Munchausen RL DQN and IQN
Lee et al., (2020) Network Randomization Proximal Policy Optimization (PPO)
Amit et al., (2020) Discount Regularization Twin Delayed DDPG (TD3)
Agarwal et al., 2021a PSM Data Regularized-Q (DrQ)
Liu et al., (2021) BN and dropout and $L_{2}/L_{1}$ PPO, TRPO, SAC, A2C
@@
Some work employs contrastive representation learning to learn deep reinforcement learning policies from state observations that are close to each other (Agarwal et al., 2021a, ). This study leverage the temporal aspect of reinforcement learning and propose a policy similarity metric. The main goal of the paper is to lay out the sequential structure and utilize representation learning to learn generalizable abstractions from state representations. One drawback of this study is that most of the experimental study is conducted in a non-baseline environment (i.e. Rectangle game and Distracting DM Control Suite). Malik et al., (2021) studies query complexity of reinforcement learning policies that can generalize to multiple environments. The authors of this study focus on an example of the transition probability transformation setting $\mathcal{G}_{\mathcal{P}}$ in Definition 3.6, and the reward function transformation setting $\mathcal{G}_{R}$ in Definition 3.4.
@@
Another line of study in direct function generalization investigates the relationship between reduced discount factor and adding an $\ell_{2}$-regularization term to the loss function, i.e. weight decay (Amit et al.,, 2020). The authors in this work demonstrate the explicit connection between reducing the discount factor and adding an $\ell_{2}$-regularizer to the value function for temporal difference learning. In particular, this study demonstrates that adding an $\ell_{2}$-regularization term to the loss function is equal to training with a lower discount term, which the authors refer to as discount regularization. The results of this study however are based on experiments from tabular reinforcement learning, and the low dimensional setting of the MuJoCo environment (Todorov et al.,, 2012). This study is also another clear example of algorithmic generalization $\mathcal{G}_{\mathbb{A}}$ as described in Definition 3.3.
@@
On the reward transformation for generalization setting $\mathcal{G}_{R}$ defined in Definition 3.4, Vieillard et al., 2020b adds the scaled log policy to the current rewards. To overcome overfitting some work tries to learn explicit or implicit similarity between the states to obtain a reasonable policy (Lan et al.,, 2021). In particular, the authors in this work try to unify the state space representations by providing a taxonomy of metrics in reinforcement learning. Several studies proposed different ways to include Kullback-Leibler divergence between the current policy and the pre-updated policy to add as a regularization term in the reinforcement learning objective (Schulman et al.,, 2015). Recently, some studies argued that utilizing Kullback-Leibler regularization implicitly averages the state-action value estimates (Vieillard et al., 2020a, ).
Table 5: Environment and algorithm details for adversarial policy regularization and attack techniques in deep reinforcement learning. Note that most of the methods based on adversarial approaches are a form of generalization assessment through state observations $\mathcal{G}^{S}$ as described in Definition 3.8, and some falls under the generalization through environment dynamics $\mathcal{G}_{\mathcal{P}}$ as described in Definition 3.6.
Citation Method Environment Algorithm
Huang et al., (2017) Fast Gradient Sign (FGSM) ALE DQN, TRPO, A3C
Kos and Song, (2017) Fast Gradient Sign (FGSM) ALE DQN & IQN
Korkmaz, (2022) Adversarial Framework ALE DDQN & A3C
Lin et al., (2017) Timing ALE A3C & DQN
Pinto et al., (2017) Zero-sum game MuJoCo RARL
Gleave et al., (2020) Adversarial Policies MuJoCo PPO
Korkmaz, (2023) Natural Attacks ALE DDQN & A3C
Huan et al., (2020) State Adversarial-DQN ALE and $L_{\textrm{M}}$ [^4] Low dimensional state MuJoCo refers to the setting of MuJoCo where the state dimensions are not represented by pixels and dimensions of the state observations range from 11 to 117. DDQN & PPO
Korkmaz, 2024c Diagnostic Adversarial Volatility ALE DDQN